OpenStudy (anonymous):

A snail travels at a rate of 2.58 feet per minute. Write a rule to describe the function. How far will the snail travel in 9 minutes?

OpenStudy (anonymous):

well it all depends on the rule divide 9 by 2.58 and make a true statement

OpenStudy (anonymous):

distance = speed * time

OpenStudy (anonymous):

2.58 * 9 = 23.22 ft

OpenStudy (anonymous):

y=total distance, x=time y=2.58x when x=9 y=2.58*9 so y=23.22ft